How many moles of PCl5 can be produced from 50.0 g of Cl2 (and excess P4)?


1
Expert's answer
2021-09-22T00:27:29-0400

The balanced chemical equation is

P4 + 10Cl2 = 4PCl5[sub]

The mass of chlorine is m(Cl[sub]2) = 59.0 g

The amount of chlorine is n(Cl2) = m(Cl2)"\/" M(Cl2) = 59"\/" 70.906 = 0.832 mol


10 moles of Cl2 yield 4 moles of PCl5;

0.832 of Cl2 yield "\times" moles of PCl5

n(PCl5) = 4"\times" 0.832"\/" 10 = 0.333 mol


Answer: 0.333 mol